OpenStudy (anonymous):

can you please solve: obtain the equation of the tangent to the curve at YX^2=1, at point A(p,1/p2) if this tangent line cutting the x-axis at point B, and y-axis at point C. PROVE that CA=2AB.

myininaya (myininaya):

that is p_2 not 2p, right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

So CA means the dot product of C and A? What kinda product is that? Also I suppose you know to find the x-intercept and y-intercept of the tangent line?

OpenStudy (anonymous):

yes but CA is where C is a point at the y-axis and A is a point at the curve of the tangent, i did't get you about what's the product? no, how?

myininaya (myininaya):

I don't know what CA=2AB means? I know what everything else means. But C, A and B are points.

myininaya (myininaya):

I can help you with everything before CA=2AB but then I can't go further unless I know the notation of what CA=2AB means. Did you find y' yet?

OpenStudy (anonymous):

i think we have to obtain the area of the tangent if i'm not mistaken. that's what CA=2AB maybe no i didn't find y' yet, if you can please solve it

myininaya (myininaya):

You have yx^2=1 You could solve for y first. Then find the derivate.

myininaya (myininaya):

\[y=\frac{1}{x^2}\] Take derivative of both sides.

OpenStudy (anonymous):

y2x=0 d/dx(yx^2)=d/dx(1) y2x=0 2xy=0 so?

myininaya (myininaya):

\[(yx^2)'=y(x^2)'+(y)'x^2 \text{ use product rule} \]

OpenStudy (anonymous):

ok, and then...

myininaya (myininaya):

Find y'.

myininaya (myininaya):

To find the slope of the tangent line going through (p,1/p2) evaluate y' at that point.

OpenStudy (anonymous):

thanks myininaya for making things clear, even though i didn't solve it because i still have another thing to do.. i really appreciate your help. Thank you

myininaya (myininaya):

Ok so you know to find the equation of the tangent line. Find the x and y intercept of that equation. You should be able to show that equality whatever that equality means because I still have no clue what CA=2AB means.

OpenStudy (anonymous):

yeah, i'll do my best to finish it.. me too i don't know

OpenStudy (vivek3461):

 Refer the drawing for the question asked. The curve is not exactly Y.X^2 = 1.

OpenStudy (vivek3461):

Take a point A (p, 1/p^2) and find the equation of tangent for the given curve at that point. Find y intercept, x intercept and then find the distance between x intercept and the tangent point (AB) and find the distance between Y intercept and the tangent point ( AC ). Then prove AB = 2 x AC

myininaya (myininaya):

Oh so AB is the distance of from point A to B. We can't do your work for you.

myininaya (myininaya):

This problem won't take very long since we know what AC=2*AB means now.

OpenStudy (anonymous):

hhhh, actually thats what i said at the begining "we have to find the area"

myininaya (myininaya):

He said distance, not area.

OpenStudy (anonymous):

almost the same

myininaya (myininaya):

Entirely different.